# A Hybrid Approach to Mitigate Errors in Linear Photonic Bell-State Measurement for Quantum Interconnects

### v1

@@ Abstract
Optical quantum information processing critically relies on Bell-state measurement, a ubiquitous operation for quantum communication and computing. Its practical realization involves the interference of optical modes and the detection of a single photon in an indistinguishable manner. Yet, in the absence of efficient photon-number resolution capabilities, errors arise from multi-photon components, decreasing the overall process fidelity. Here, we introduce a novel hybrid detection scheme for Bell-state measurement, leveraging both on-off single-photon detection and quadrature conditioning via homodyne detection. We derive explicit fidelities for quantum teleportation and entanglement swapping processes employing this strategy, demonstrating its efficacy. We also compare with photon-number resolving detectors and find a strong advantage of the hybrid scheme in a wide range of parameters. This work provides a new tool for linear optics schemes, with applications to quantum state engineering and quantum interconnects.
@@ II.1 POVM derivation
The description of this POVM element relies on the matrices of the beamsplitter operation $\mathrm{BS}(R)$ that depends on its reflectivity, the single-photon detection POVM element $\Pi^{\mathrm{on}}_{\mathrm{SPD}}$, which corresponds here to an on-off detection [27], and the homodyne detection POVM elements [28], integrated over the detection window $\Delta$ and independent of the detection phase $\theta$, i.e.,
@@
Figure 2: Diagonal elements of the HBSM POVM $\Pi^{\textrm{on}}$ element. They are given for a fixed reflectivity $R=0.1$ and for different conditioning windows $\Delta$, from a large acceptance to a very small one relative to shot noise: (a) $\Delta=10$, (b) $\Delta=1$, (c) $\Delta=0.25$ and (d) $\Delta=0.001$. The coloring indicates the wanted green $\Pi^{\mathrm{on}}_{[1,1]}$ and unwanted red $\Pi^{\mathrm{on}}_{[2,2]}$ elements. All matrices were normalized for comparability, with normalization factors $\mathcal{N}_{a}=0.824$, $\mathcal{N}_{b}=0.156$, $\mathcal{N}_{c}=0.037$ and $\mathcal{N}_{d}=0.00015$. The detection efficiencies are set to $\eta_{\mathrm{HD}}~{}=~{}\eta_{\mathrm{SPD}}~{}=~{}0.9$.
As this measurement is phase-independent, the POVM element has only diagonal terms that are non-zero. As mentioned in the introduction we mostly assume components in the Fock basis up to ${\left|{2}\right\rangle}$. The diagonal elements are given by
@@
where $\eta_{\mathrm{SPD}}$ is the detection efficiency of the single-photon detector and $\eta_{\mathrm{HD}}$ the efficiency of the homodyne detection. Many implementations require a filtering stage before the single-photon detection, and its efficiency can be integrated into an overall $\eta_{\mathrm{SPD}}$. In the asymptotic limit of small $R$ and $\Delta$, the POVM elements reduce to
@@
As the ideal HBSM shall project on the single-photon state, we aim at reducing $\Pi^{\mathrm{on}}_{[2,2]}$ while maximizing $\Pi^{\mathrm{on}}_{[1,1]}$ with the tunable parameters $R$ and $\Delta$. As shown in Fig. 2, for a fixed reflectivity $R$ this is obtained by decreasing the conditioning window $\Delta$. From a certain value of $\Delta$, typically 0.25, decreasing it further has a limited effect as can be seen in Fig. 2(c) and 2(d). Furthermore, we note that we are actually achieving a measurement that detects odd Fock states with higher probability than even ones, as expected from quadrature conditioning at the origin. In our case of interest for BSM, this feature reduces to a boosted single-photon versus two-photon detection. To further explore the HBSM in this case a benchmark needs to be defined, as done in the following.
@@ II.2 Projectivity benchmark: purity of the HBSM
The ideal HBSM should act as a perfect single-photon detector, having only one non-zero element, i.e., $\Pi^{\textrm{on}}_{[1,1]}~{}=~{}1$. In order to benchmark the HBSM we propose the purity $\mathcal{P}(\Pi^{\textrm{on}})$. As above, the POVM will be truncated such that we only consider the first diagonal elements defined in Eq. II.1. The purity can then be defined as
@@
and only depends on the homodyne measurement efficiency $\eta_{\mathrm{HD}}$. For $\eta_{\mathrm{HD}}\rightarrow 1$, the purity reaches unity, corresponding to a projective measurement.
Given this expression, we can now evaluate how the purity depends on $\Delta$ and $R$. This is detailed in Fig. 3(a) and 3(c), respectively, and for two values of SPD detection efficiency $\eta_{\mathrm{SPD}}$. Figures 3(b) and 3(d) provide the maximal possible detection probability $P_{\mathrm{max}}^{\,\mathrm{on}}$, which is defined as the maximal detection probability over all states, thereby providing the probability for the most favorable state. The reflectivity $R$ has a larger impact on the purity while $\Delta$ shows a more favorable trade-off between purity and detection probability. We note also that a decrease in the SPD efficiency increases the importance of the two-photon component.
@@ III Application to the hybrid Bell-state Measurement
Figure 4: Effect of the HBSM applied to one mode of a two-mode state with photon-number correlations, of the form $\sqrt{c_{11}}{\left|{11}\right\rangle}+\sqrt{1-c_{11}}{\left|{22}\right\rangle}$ or $c_{11}{\left|{11}\right\rangle}{\left\langle{11}\right|}+(1-c_{11}){\left|{22}\right\rangle}{\left\langle{22}\right|}$ where $c_{11}$ is the single-photon weight. The discrepancy from the ideal case, in which the HBSM projects the second mode onto ${\left|{1}\right\rangle}$, is given by the fidelity $\mathcal{F}$ as a function of the weight $c_{11}$. The fidelity is displayed for (a) different reflectivities $R$ with fixed conditioning window $\Delta=0.1$ and for (c) different conditioning windows $\Delta$ with fixed $R=0.1$. The black dotted lines indicate the fidelity without any measurement. (b) and (d) provide the corresponding success probabilities $P^{\mathrm{on}}$. The detection efficiencies are set to $\eta_{\mathrm{HD}}=0.9$ and $\eta_{\mathrm{SPD}}=0.5$.
@@ III.1 The HBSM as a single-photon detector
Figure 6: Teleportation protocol and hybrid Bell-state measurement. (a) A qubit to be teleported ${\left|{\psi}\right\rangle}_{in}=\sqrt{c_{0}}{\left|{0}\right\rangle}+\sqrt{c_{1}}{\left|{1}\right\rangle}$ undergoes a BSM with one mode of a Bell state. The remaining mode is ideally projected onto the input state upon successful detection. (b) The fidelity $\mathcal{F}_{\mathrm{T}}$ between the input and teleported states and (c) the success probability $P^{\mathrm{on}}$ are given as a function of the input single-photon weight $c_{1}$ for different reflectivities $R$ of the HBSM. The black dashed line indicates the worst case scenario, i.e., using an on-off detector with perfect efficiency. The conditioning window is fixed at $\Delta=0.1$ and the detection efficiencies are set to $\eta_{\mathrm{HD}}=0.9$ and $\eta_{\mathrm{SPD}}=0.5$.
@@ III.2 The HBSM in a teleportation protocol
Quantum teleportation as shown in Fig. 6(a) considers an input qubit of the form ${\left|{\varphi}\right\rangle}_{\mathrm{in}}~{}=~{}\sqrt{c_{0}}~{}{\left|{0}\right\rangle}~{}+~{}\sqrt{c_{1}}{\left|{1}\right\rangle}$ and one of the four entangled Bell states ${\left|{\phi}\right\rangle}^{\pm}~{}=~{}(1/\sqrt{2})({\left|{00}\right\rangle}\pm{\left|{11}\right\rangle})$ or ${\left|{\psi}\right\rangle}^{\pm}=(1/\sqrt{2})({\left|{01}\right\rangle}\pm{\left|{10}\right\rangle})$ as a resource. A Bell-state measurement between one mode of the Bell state and the input state will then project the remaining mode into the input state up to a local unitary, as reminded here [29]:
${\left|{\varphi}\right\rangle}_{\mathrm{in}}\otimes{\left|{\psi}\right\rangle}^{+}\propto\;$ ${\left|{\psi}\right\rangle}^{+}\otimes$ ${\left|{\varphi}\right\rangle}_{\mathrm{in}}$ $+$ ${\left|{\psi}\right\rangle}^{-}\otimes$ $Z{\left|{\varphi}\right\rangle}_{\mathrm{in}}+$
${\left|{\phi}\right\rangle}^{+}\otimes$ $X{\left|{\varphi}\right\rangle}_{\mathrm{in}}$ $+$ ${\left|{\phi}\right\rangle}^{-}\otimes$ $ZX{\left|{\varphi}\right\rangle}_{\mathrm{in}}.$ (7)
@@
[…]
Only the states ${\left|{\psi}\right\rangle}^{\pm}$ can be detected with this scheme and this detection relies on the fact that one can detect exactly a single photon. For simplification, we will only consider ${\left|{\psi}\right\rangle}^{+}$. The probability to measure a single photon is $0.25$, while the probability to measure two photons is $c_{1}/4$. In the usual BSM scheme, an on-off detector is used. As it detects single-photon and two-photon components equally, we can define the worst case scenario of the teleportation fidelity $\mathcal{F}_{\mathrm{T}}$ by using an on-off detector with perfect efficiency.
@@ III.3 The HBSM in a swapping protocol
where the BSM is applied between each second mode. After one mode of each interacted on a balanced beamsplitter, a projection onto a single-photon state will lead to a Bell state ${\left|{\psi}\right\rangle}^{\pm}$, dependent on the detector. We will again consider the detection on only one output port, projecting the final modes into $(1/\sqrt{2})({\left|{01}\right\rangle}+{\left|{10}\right\rangle})$. Equal detection of single photons with probability $c_{10}(1-c_{10})$ and two photons $0.5(1-c_{10})^{2}$ leads to the worst case of the swapping fidelity $\mathcal{F}_{\mathrm{S}}$ between the ideal output state ${\left|{\psi^{\textrm{out}}}\right\rangle}$ and the actual output state, defined by the use of an ideal on-off detector.
@@ IV Benchmarking the HBSM against photon-number resolving detectors
Figure 10: Comparison of BSM implementations for quantum teleportation and entanglement swapping. (a) and (b) provide the fidelity $\mathcal{F}_{\mathrm{T}}$ and success probability $P^{\,\mathrm{on}}$ for the teleportation of a balanced qubit. (c) and (d) provide the fidelity $\mathcal{F}_{\mathrm{S}}$ and success probability $P^{\,\mathrm{on}}$ for the swapping of two balanced ${\left|{\phi^{+}}\right\rangle}$ states. Two values of conditioning window $\Delta$ are considered for HBSM, and compared to implementations with $N=1$, $N=2$, and $N=\infty$ that emulates perfect photon-number resolution. All POVM elements are truncated at $n=2$. For HBSM, the homodyne detection efficiency is set to $\eta_{\mathrm{HD}}=0.9$, and the reflectivity is fixed at $R=0.1$.

### v2

@@ Abstract
Optical quantum information processing critically relies on Bell-state measurement, a ubiquitous operation for quantum communication and computing. Its practical realization involves the interference of optical modes and the detection of a single photon in an indistinguishable manner. Yet, in the absence of efficient photon-number resolution capabilities, errors arise from multi-photon components, decreasing the overall process fidelity. Here, we introduce a hybrid detection scheme for Bell-state measurement, leveraging both on-off single-photon detection and quadrature conditioning via homodyne detection. We derive explicit fidelities for quantum teleportation and entanglement swapping processes employing this strategy, demonstrating its efficacy. We also compare with photon-number resolving detectors and find a strong advantage of the hybrid scheme in a wide range of parameters. This work provides a new tool for linear optics schemes, with applications to quantum state engineering and quantum interconnects.
@@ II.1 POVM derivation
The description of this POVM element relies on the matrices of the beamsplitter operation $\mathrm{BS}(R)$ that depends on its reflectivity, the single-photon detection POVM element $\Pi^{\mathrm{on}}_{\mathrm{SPD}}$, which corresponds here to an on-off detection [27], and the homodyne detection POVM elements [28] (see Appendix A), integrated over the detection window $\Delta$ and independent of the detection phase $\theta$, i.e.,
@@
Figure 2: Diagonal elements of the HBSM POVM $\Pi^{\textrm{on}}$ element. They are given for a fixed reflectivity $R=0.1$ and for different conditioning windows $\Delta$, from a large acceptance to a very small one relative to shot noise: (a) $\Delta=10$, (b) $\Delta=1$, (c) $\Delta=0.25$ and (d) $\Delta=0.001$. The coloring indicates the wanted green $\Pi^{\mathrm{on}}_{[1,1]}$ and unwanted red $\Pi^{\mathrm{on}}_{[2,2]}$ elements. All matrices were normalized for comparability, with normalization factors $\mathcal{N}_{a}=0.824$, $\mathcal{N}_{b}=0.156$, $\mathcal{N}_{c}=0.037$ and $\mathcal{N}_{d}=0.00015$. The detection efficiencies are set to $\eta_{\mathrm{HD}}~{}=~{}\eta_{\mathrm{SPD}}~{}=~{}0.9$. Diagonal elements $n>2$ are indicated in grey. In our study, their contributions will be negligible because the states to be measured have a very low probability of carrying more than two photons.
As this measurement is phase-independent, the POVM element has only diagonal terms that are non-zero. The full expressions of the first elements are given in Appendix B. As mentioned in the introduction we mostly assume components in the Fock basis up to ${\left|{2}\right\rangle}$. The diagonal elements of interest can be read as
@@
where $\eta_{\mathrm{SPD}}$ is the detection efficiency of the single-photon detector and $\eta_{\mathrm{HD}}$ the efficiency of the homodyne detection. Many implementations require a filtering stage before the single-photon detection, and its efficiency can be integrated into an overall $\eta_{\mathrm{SPD}}$.
As the ideal HBSM shall project on the single-photon state, we aim at reducing $\Pi^{\mathrm{on}}_{[2,2]}$ while maximizing $\Pi^{\mathrm{on}}_{[1,1]}$ with the tunable parameters $R$ and $\Delta$. As shown in Fig. 2, for a fixed reflectivity $R$ this is obtained by decreasing the conditioning window $\Delta$. From a certain value of $\Delta$, typically 0.25, decreasing it further has a limited effect as can be seen in Fig. 2(c) and 2(d). Furthermore, we note that we are actually achieving a measurement that detects odd Fock states with higher probability than even ones, as expected from quadrature conditioning at the origin. In our case of interest for BSM, this feature reduces to a boosted single-photon versus two-photon detection, which can be more easily seen in the asymptotic limit of small $R$ and $\Delta$, where the POVM elements reduce to
@@
To further explore the HBSM in this case a benchmark needs to be defined, as done in the following.
@@ II.2 Projectivity benchmark: purity of the HBSM
The ideal HBSM should act as a perfect single-photon detector, having only one non-zero element, i.e., $\Pi^{\textrm{on}}_{[1,1]}~{}=~{}1$. In order to benchmark the HBSM we propose the purity $\mathcal{P}(\Pi^{\textrm{on}})$, indicating how close the HBSM is to an ideal projective measurement. The POVM will be truncated such that we only consider the first diagonal elements defined in Eq. II.1. As above, this is due to the small higher order Fock components in the considered input states. The purity can then be defined as
@@
and only depends on the homodyne measurement efficiency $\eta_{\mathrm{HD}}$. For $\eta_{\mathrm{HD}}\rightarrow 1$, the purity reaches unity, corresponding to a projective measurement with $\Pi^{\textrm{on}}_{[1,1]}=1$.
Given this expression, we can now evaluate how the purity depends on $\Delta$ and $R$. This is detailed in Fig. 3(a) and 3(c), respectively, and for two values of SPD detection efficiency $\eta_{\mathrm{SPD}}$. Figures 3(b) and 3(d) provide the maximal possible detection probability $P_{\mathrm{max}}^{\,\mathrm{on}}$, which is defined as the maximal detection probability over all states $P_{\mathrm{max}}^{\,\mathrm{on}}=\max_{\rho}\mathrm{Tr}[\Pi^{\mathrm{on}}\rho]$, thereby providing the detection probability for the most favorable state. The reflectivity $R$ has a larger impact on the purity while $\Delta$ shows a more favorable trade-off between purity and detection probability. We note also that a decrease in the SPD efficiency increases the importance of the two-photon component.
@@ III Application to the hybrid Bell-state Measurement
Figure 4: Effect of the HBSM applied to one mode of a two-mode state with photon-number correlations, of the form $\sqrt{c_{11}}{\left|{11}\right\rangle}+\sqrt{1-c_{11}}{\left|{22}\right\rangle}$ or $c_{11}{\left|{11}\right\rangle}{\left\langle{11}\right|}+(1-c_{11}){\left|{22}\right\rangle}{\left\langle{22}\right|}$ where $c_{11}$ is the single-photon weight. The discrepancy from the ideal case, in which the HBSM projects the second mode onto ${\left|{1}\right\rangle}$, is given by the fidelity $\mathcal{F}$ as a function of the weight $c_{11}$. The fidelity is displayed for (a) different reflectivities $R$ with fixed conditioning window $\Delta=0.1$ and for (c) different conditioning windows $\Delta$ with fixed $R=0.1$. The black dotted lines indicate the fidelity without any measurement. (b) and (d) provide the corresponding success probabilities $P^{\mathrm{on}}$. For the blue case in (d), the probability for $c_{11}=0.5$ is equal to $2.10^{-3}$. The detection efficiencies are set to $\eta_{\mathrm{HD}}=0.9$ and $\eta_{\mathrm{SPD}}=0.5$.
@@ III.1 The HBSM as a single-photon detector
Figure 6: Teleportation protocol and hybrid Bell-state measurement. (a) A qubit to be teleported ${\left|{\varphi}\right\rangle}_{in}=\sqrt{c_{0}}{\left|{0}\right\rangle}+\sqrt{c_{1}}{\left|{1}\right\rangle}$ undergoes a BSM with one mode of a Bell state. The remaining mode is ideally projected onto the input state upon successful detection. (b) The fidelity $\mathcal{F}_{\mathrm{T}}$ between the input and teleported states and (c) the success probability $P^{\mathrm{on}}$ are given as a function of the input single-photon weight $c_{1}$ for different reflectivities $R$ of the HBSM. The black dashed line indicates the SPD case, i.e., using an on-off detector with perfect efficiency, and without homodyning. The conditioning window is fixed at $\Delta=0.1$ and the detection efficiencies are set to $\eta_{\mathrm{HD}}=0.9$ and $\eta_{\mathrm{SPD}}=0.5$.
@@ III.2 The HBSM in a teleportation protocol
Quantum teleportation as shown in Fig. 6(a) considers an input qubit of the form ${\left|{\varphi}\right\rangle}_{\mathrm{in}}~{}=~{}\sqrt{c_{0}}~{}{\left|{0}\right\rangle}~{}+~{}\sqrt{c_{1}}{\left|{1}\right\rangle}$ and one of the four entangled Bell states ${\left|{\phi}\right\rangle}^{\pm}~{}=~{}(1/\sqrt{2})({\left|{00}\right\rangle}\pm{\left|{11}\right\rangle})$ or ${\left|{\psi}\right\rangle}^{\pm}=(1/\sqrt{2})({\left|{01}\right\rangle}\pm{\left|{10}\right\rangle})$ in the photon-number basis as a resource. A Bell-state measurement between one mode of the Bell state and the input state will then project the remaining mode into the input state up to a local unitary, as reminded here [29]:
${\left|{\varphi}\right\rangle}_{\mathrm{in}}\otimes{\left|{\psi}\right\rangle}^{+}\propto\;$ ${\left|{\psi}\right\rangle}^{+}\otimes$ ${\left|{\varphi}\right\rangle}_{\mathrm{out}}$ $+$ ${\left|{\psi}\right\rangle}^{-}\otimes$ $Z{\left|{\varphi}\right\rangle}_{\mathrm{out}}+$
${\left|{\phi}\right\rangle}^{+}\otimes$ $X{\left|{\varphi}\right\rangle}_{\mathrm{out}}$ $+$ ${\left|{\phi}\right\rangle}^{-}\otimes$ $ZX{\left|{\varphi}\right\rangle}_{\mathrm{out}}.$ (7)
@@
$\mathrm{BS}(0.5){\left|{\psi}\right\rangle}^{+}$ $=$ $\;{\left|{01}\right\rangle}$
$\mathrm{BS}(0.5){\left|{\psi}\right\rangle}^{-}$ $=$ $\;{\left|{10}\right\rangle}$
$\mathrm{BS}(0.5){\left|{\phi}\right\rangle}^{\pm}$ $=$ $\;\frac{1}{\sqrt{2}}{\left|{00}\right\rangle}\pm\frac{1}{2}({\left|{02}\right\rangle}-{\left|{20}\right\rangle}),$ (8)
Only the states ${\left|{\psi}\right\rangle}^{\pm}$ can be detected with this scheme and this detection relies on the fact that one can detect exactly a single photon. For simplification, we will only consider ${\left|{\psi}\right\rangle}^{+}$. The probability to measure a single photon is $0.25$, while the probability to measure two photons is $c_{1}/4$, and therefore depends on the single-photon weight $c_{1}$ of the input qubit. In the usual BSM scheme, an on-off detector is used. As it detects single-photon and two-photon components equally, we can define the worst case scenario of the teleportation fidelity $\mathcal{F}_{\mathrm{T}}$ by using an on-off detector with perfect efficiency.
@@ III.3 The HBSM in a swapping protocol
where the BSM is applied between each second mode. After one mode of each interacted on a balanced beamsplitter, a projection onto a single-photon state will lead to a Bell state ${\left|{\psi}\right\rangle}^{\pm}$, dependent on the detector. We will again consider the detection on only one output port, projecting the final modes into $(1/\sqrt{2})({\left|{01}\right\rangle}+{\left|{10}\right\rangle})$. Equal detection of single photons with probability $c_{01}(1-c_{01})$ and two photons $0.5c_{01}^{2}$ leads to the worst case of the swapping fidelity $\mathcal{F}_{\mathrm{S}}$ between the ideal output state ${\left|{\psi^{\textrm{out}}}\right\rangle}$ and the actual output state, defined by the use of an ideal on-off detector.
@@ IV Benchmarking the HBSM against photon-number resolving detectors
Figure 10: Comparison of BSM implementations for quantum teleportation and entanglement swapping. (a) and (b) provide the fidelity $\mathcal{F}_{\mathrm{T}}$ and success probability $P^{\,\mathrm{on}}$ for the teleportation of a balanced qubit. (c) and (d) provide the fidelity $\mathcal{F}_{\mathrm{S}}$ and success probability $P^{\,\mathrm{on}}$ for the swapping of two balanced ${\left|{\phi^{+}}\right\rangle}$ states. Two values of conditioning window $\Delta$ are considered for HBSM, and compared to implementations with $N=1$, $N=2$, $N=5$, and $N=\infty$ that emulates perfect photon-number resolution. All POVM elements are truncated at $n=2$. For HBSM, the homodyne detection efficiency is set to $\eta_{\mathrm{HD}}=0.9$, and the reflectivity is fixed at $R=0.1$.
@@
Indeed PNR detectors are not largely commercially available at the moment, and the threshold value we obtain is comparable to the best values demonstrated for PNR detectors, reachable at a limited number of wavelengths.
@@ Acknowledgements.
Appendix A Definition of the homodyne POVM
As used in our HBSM POVM derivation, here we remind the POVM element corresponding to a homodyne detection [28]
$\text{HD}(x,\theta)$ $=$ ${\left|{x,\theta}\right\rangle}{\left\langle{x,\theta}\right|}$
${\left|{x,\theta}\right\rangle}$ $=$ ${\langle{x_{\theta}|n}\rangle}$
[…]
which is the wavefunction of Fock states projected onto the quadrature $x_{\theta}$. We can account for an imperfect detection efficiency $\eta$, by applying a beamsplitter operation, called Bernoulli transformation $B$. The resulting homodyne POVM elements can be written as
$\Pi_{\eta}(x,\theta)$ $=$ $\sum_{m,n,i}B_{m+i,m}(\eta)\;B_{n+i,n}(\eta)\;$
$\qquad\qquad\text{HD}(x,\theta){\left|{n+i}\right\rangle}{\left\langle{m+i}\right|}$
$B_{m+i,m}$ $=$ $\sqrt{\binom{m+i}{m}\eta^{m}(1-\eta)^{i}}.$
We can therefore interpret this measurement as sampling from the marginal distribution along $x_{\phi}$ of the Wigner function.
Appendix B HBSM POVM elements
We provide the expression for the HBSM POVM elements up to $n=4$ as displayed in Fig. 2.
$\Pi^{\mathrm{on}}_{[0,0]}$ $=$ 00 0
$\Pi^{\mathrm{on}}_{[2,2]}$ $=$ $R\;\eta_{\mathrm{SPD}}\Big{[}\text{erf}\left(\frac{\Delta}{2}\right)(2-R\;\eta_{\mathrm{SPD}})$
$\Pi^{\mathrm{on}}_{[3,3]}$ $=$ $R\;\eta_{\mathrm{SPD}}\;\text{erf}\left(\frac{\Delta}{2}\right)(\eta_{\mathrm{SPD}}\;R(\eta_{\mathrm{SPD}}\;R-3)+3)$
$-\frac{1}{4\sqrt{\pi}}3\;e^{-\frac{\Delta^{2}}{4}}\Delta\;\eta_{\mathrm{SPD}}\eta_{\mathrm{HD}}(R-1)R$
×$\times\Big{[}\left(\Delta^{2}-6\right)\eta_{\mathrm{HD}}(R-1)+4\;\eta_{\mathrm{SPD}}\;R-8\Big{]}$
$\Pi^{\mathrm{on}}_{[4,4]}$ $=$ $\frac{1}{6}\eta_{\mathrm{SPD}}\;R\Biggl{(}\frac{1}{\sqrt{\pi}}e^{-\frac{\Delta^{2}}{4}}\Delta\;\eta_{\mathrm{HD}}(R-1)\Big{[}9\left(\Delta^{2}-6\right)$
×$\times\eta_{\mathrm{HD}}(R-1)(\eta_{\mathrm{SPD}}\;R-2)+\left(\Delta^{4}-20\Delta^{2}+60\right)$
×$\times\eta_{\mathrm{HD}}^{2}(R-1)^{2}+24\;\eta_{\mathrm{SPD}}\;R(\eta_{\mathrm{SPD}\;}R-3)+72\Big{]}$
$-6\;\text{erf}\left(\frac{\Delta}{2}\right)(\eta_{\mathrm{SPD}}\;R-2)$
×$\times(\eta_{\mathrm{SPD}}\;R(\eta_{\mathrm{SPD}}\;R-2)+2)\Biggr{)}.$